OpenStudy (anonymous):

The average of Jim's two test scores must be at least 90 to make an A in the class. Jim got a 95 on his first test. What grades can Jim get on his second test to make an A in the class?

OpenStudy (phi):

do you know the rule for finding the average of 2 numbers?

OpenStudy (phi):

are you saying you don't know how to find the average of (for example) 3 and 5 ?

OpenStudy (mathmale):

PP: I'd suggest you look up "average" and/or "mean" and inform yourself by doing that what an "average" is and how one would go about finding it. You will run into the concept of "average" often if you continue to take math or math-related course work.

OpenStudy (amtran_bus):

Average is the sum of the numbers divided by however many numbers there were. Therefore, the average of 3 and 5 are 3+5=7 Since there were 2 numbers, 7/2 =3.5

OpenStudy (amtran_bus):

So, Jim to make a 90 or above in his class, he needs to make nothing less than an 85 on his next test. Proof: 85+95=180 180/2=90 84+95=179 179/2=89.5 and trust me some professors WILL NOT ROUND UP! lol

OpenStudy (anonymous):

Ted needs an average of at least 70 on his three history tests. He has already scored 85 and 60 on two tests. What is the minimum grade Ted needs on his third test?

OpenStudy (amtran_bus):

I worked one, you can do this one.

OpenStudy (mathmale):

Let x represent the third test score. Ted has already scored 85 and 60 on two previous tests. Using that variable ' x ' and the other two test scores, find the sum of the three test scores (in other words, write an expression, because you don't yet know the value of x).